OpenStudy (anonymous):

Molly exercised for 15 minutes on monday and 15 minutes on tuesday. she spent 2/3 of her excercise time running. How many minutes did molly run on monday and tuesday combined?

OpenStudy (amistre64):

15+15 = 30 minutes total time for mon and tues. 2/3 of that time was running.. 1/3 of 30 = 10; so 2/3 of 30 = 20 20 minutes running total